# textwidth=0.8textheight=0.9

### v1

@@ 4.1 Architecture and Training Strategies
To achieve this objective, the network parameters are iteratively updated using backpropagation [193, 194, 195]; a procedure that computes the gradients of the loss function with respect to all trainable parameters. These gradients provide the necessary direction for optimization algorithms to adjust the network parameters toward minimizing the loss. Commonly used first-order optimization methods include stochastic gradient descent (SGD) [196, 197] and the Adam optimizer [198], both of which rely on gradient information to efficiently update the model during training. In contrast, second-order or quasi-Newton methods, such as the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [199] and its limited-memory variant (L-BFGS) [200], utilize curvature information to accelerate convergence. More recent advancements, such as the Self-Scaled Broyden (SSB) algorithms [201, 107], further enhance these quasi-Newton approaches by employing self-scaling strategies. These quasi-Newton methods iteratively construct approximations to the inverse Hessian matrix, allowing for more informed parameter updates by incorporating curvature information. This typically results in faster and more stable convergence compared to first-order methods, especially in high-dimensional or ill-conditioned optimization problems [202, 203]. To address the optimization challenges in data-driven KANs, it is worth considering hybrid strategies that combine the strengths of different optimizers. For instance, combining adaptive first-order methods like Adam [198], which are effective for fast initial convergence, with second-order methods such as L-BFGS [204, 200], which offer precise updates near local optima, may significantly improve training efficiency. This approach can be particularly beneficial for KANs, where the optimization involves a large number of parameters and the joint tuning of both the shape and location of spline-based transformations [166, 167, 148]. In this context, Adam can help navigate the early stages of training with noisy gradients, while L-BFGS can refine the solution more effectively in the later stages.
@@
Recent advances in the design of data-driven KAN architectures have centered on structural innovations aimed at improving expressivity, interpretability, and efficiency. One direction involves hybrid basis function integration, as demonstrated in BSRBF-KAN, which combines B-splines with radial basis functions to enhance local approximation accuracy [205]. This hybrid formulation is implemented through a combined transformation layer applied to normalized inputs. The authors highlighted that, while both basis types contribute to performance, ablation studies suggest that B-splines play a more dominant role in preserving generalization. Additionally, Ta [205] provided insights into model stability and convergence behavior, indicating BSRBF-KAN’s potential as a flexible template for further exploration of basis function combinations in KANs. Other works, such as FC-KAN [192], propose function composition strategies by combining multiple elementary functions, such as wavelets, B-splines, and Gaussians, using element-wise operations or polynomial mappings to boost learning capacity. These methods show improved accuracy at the cost of increased training time, highlighting a trade-off between representational richness and computational load. Moreover, Mostajeran & Faroughi [46] proposed physics-guided parallel and serial KAN architectures with Chebyshev basis functions to evaluate their effectiveness in accurately modeling complex datasets that represent elasto-plastic material behavior. These structural adaptations reflect the growing interest in tailoring KAN frameworks to specific task demands, offering pathways toward more flexible and application-aware neural architectures.
@@
Alongside structural enrichment, another key research direction involves simplifying data-driven KAN architectures to reduce computational overhead without compromising performance. The HKAN model [206] introduces a novel hierarchical design that replaces iterative backpropagation with a non-iterative, layer-wise least squares fitting procedure, eliminating gradient computation. This architecture adopts a stacking mechanism in which each layer incrementally refines the previous output, enhancing model stability and interpretability while remaining computationally efficient. Meanwhile, the PRKAN architecture [207] targets parameter reduction by combining convolutional layers, pooling, attention mechanisms, and dimension summation strategies to drastically lower the number of learnable parameters. Despite reduced complexity, PRKAN maintains competitive performance in image classification, leveraging normalization and hybrid basis functions like GRBFs. This reflects a shift toward lightweight, scalable KAN architectures, balancing structural simplicity with functional robustness.
@@ 4.2 Applications
The adoption of data-driven KANs has been on the rise for a diverse array of tasks, including forecasting, regression, classification, and segmentation applications in both structured tabular domains and high-dimensional vision tasks. These models leverage the flexible function approximation capabilities to learn complex mappings directly from high-dimensional data. A summary of recent application-oriented advances in data-driven KANs is provided in Tables 2.
@@
EPi-cKAN Developed physics-guided serial and parallel architectures to model nonlinear elasto-plastic behavior of granular materials; Used Chebyshev polynomials as basis functions; Applied to predict stress–strain responses. [46]
@@
Time-Series KAN Developed a KAN framework for time series forecasting; Used B-spline basis functions (degree 3 and grid size 5); Applied to forecast real-world satellite traffic using normalized hourly data from the 5G-STARDUST project. [168]
@@
Geotechnical KAN Developed a KAN model to predict settlement and residual strength after liquefaction; Used B-spline basis functions; Applied to field and lab data from case studies, including risk assessment in Patna and railway embankment analysis. [86]
@@
COEFF-KANs Developed a two-stage modeling strategy to predict Coulombic Efficiency (CE) of lithium-metal battery electrolytes; Applied to a real-world Li/Cu half-cell dataset to predict log-transformed CE values. [208]
@@
SCKansformer Developed a fine-grained classification model to improve diagnostic accuracy in hematological malignancies by addressing limitations of existing microimage analysis methods; Applied to over 10,000 cell images across three datasets (BMCD-FGCD, PBC, ALL-IDB). [209]
@@
U-KAN Developed a KAN-enhanced convolutional architecture to improve nonlinear representation and interpretability in image segmentation and generative tasks; Used tokenized KAN layers; Applied to diverse imaging datasets (BUSI, GlaS, and others). [84]
@@
KACQ-DCNN Developed a classical–quantum hybrid neural architecture; Used B-spline basis functions within a dual-channel design to model nonlinear activation patterns and capture complex feature relationships; Applied to structured data for binary classification. [210]
@@
In the context of regression tasks, Vaca-Rubio et al. [168] proposed a data-driven KAN architecture as a parameter-efficient and interpretable alternative to traditional MLPs for time series forecasting, with a specific focus on satellite traffic prediction. Inspired by KART, their model replaces conventional weight-activation pairs with spline-parametrized univariate transformations, enabling greater modeling flexibility. As shown in Fig. 4(a), the KAN model closely follows the true traffic trajectory, particularly during rapid fluctuations. This demonstrates KAN’s ability to capture high-frequency temporal features and long-range dependencies. Further analysis demonstrates the impact of KAN-specific hyperparameters, such as node count and grid size, in balancing model accuracy and computational cost. These findings establish KANs as a promising, resource-efficient approach for complex time series forecasting tasks. Beyond temporal forecasting, KANs have also been successfully applied to structured regression problems. For example, Karakaş [86] employed a data-driven KAN model to predict liquefaction-induced settlements using a dataset derived from the Pohang earthquake region in South Korea, previously compiled by Park et al. [211]. The input features included depth, unit weight, cyclic stress ratio, and corrected SPT blow count, with settlement magnitude as the target variable. The study demonstrated that KANs effectively capture the nonlinear dependencies between geotechnical parameters and settlement outcomes, achieving high predictive accuracy across standard evaluation metrics such as $R^{2}$, Mean Absolute Error (MAE), and Root Mean Squared Error (RMSE). Notably, the model identified cyclic stress ratio and SPT blow count as the most influential predictors, aligning well with domain-specific understanding. These results emphasize the potential of KANs as reliable and interpretable tools for modeling complex phenomena in data-limited environments. Further extending the applicability of KANs to granular material modeling, Mostajeran & Faroughi [46] developed Chebyshev-based KAN variants to predict nonlinear elasto-plastic behavior under complex loading. Using parallel and serial KAN architectures, the models achieved accurate predictions of plastic shear strain $\gamma^{p}$ and mean effective stress $p$, even under unseen undrained loading paths such as $\xi=-4$, as shown in Fig. 4(b). These results demonstrate the effectiveness of KANs in capturing stress–strain responses in elasto-plastic materials with compact and interpretable architectures. Further broadening the landscape of KAN applications, Li et al. [208] introduced COEFF-KAN, a two-stage framework leveraging Kolmogorov–Arnold Networks to predict Coulombic Efficiency (CE) in lithium battery electrolytes. In this structured regression task, the model inputs consisted of SMILES-based molecular embeddings weighted by their molar ratios, derived from a pre-trained MoLFormer transformer. These formulation-level descriptors were then passed to a downstream KAN model, enabling the prediction of logarithmic CE (LCE) with state-of-the-art accuracy. As depicted in Fig. 4(c), the model exhibits robust generalization, with parity plots (top) demonstrating tight alignment between predicted and actual LCE values across held-out samples, and ablation studies (bottom) confirming KAN’s superior performance. The study underscores KAN’s strengths in capturing complex, high-dimensional structure–property relationships from limited chemical data, while preserving interpretability through its modular, spline-based transformations.
@@
Figure 4: (a) KAN with learnable B-spline activations for time series forecasting. The forecast is based on real GEO satellite traffic data (hourly resolution, 168h context / 24h prediction) provided by the 5G-STARDUST project, highlighting KAN’s applicability to AI-driven satellite resource management. This panel is adopted from [168]. (b) Prediction of plastic shear strain $\gamma^{p}$ and mean effective stress $p$ using parallel and serial KAN architectures under an undrained loading path with $\xi=-4$, $p_{\text{in}}=375$ kPa, and $e_{\text{in}}=0.64$. The results are reproduced based on the method presented in [46], along with the ground truth obtained from numerical integration. (c) Application of COEFF-KAN to electrolyte property prediction. Top: Parity plot comparing predicted and ground-truth values of logarithmic Coulombic Efficiency (LCE) across test samples, illustrating strong agreement and model generalization. Bottom: RMSE performance of KAN versus MLP across varying network depths and widths. This panel is reproduced from [208]. (d) Channel-wise activation maps for explainability evaluation in tumor segmentation. Integrating KAN layers enhances spatial alignment between model attention and ground truth, improving plausibility in localization tasks. This panel is adopted from [84].
@@
In the context of classification tasks, Ta [205] introduced BSRBF-KAN to evaluate two standard classification benchmarks, MNIST and Fashion-MNIST, using a consistent network structure across all experiments. BSRBF-KAN achieved stable and competitive performance, reporting average accuracies of 97.55% on MNIST and 89.33% on Fashion-MNIST, while also exhibiting faster convergence and smoother loss decay across training epochs. This study revealed the importance of architectural components such as layer normalization and base output terms, while highlighting that B-spline components contributed more significantly to performance than RBFs. These results emphasize the potential of hybrid basis-function strategies within the KAN framework for enhancing expressivity and robustness in classification tasks. In the same line of KAN-based classification work, Chen et al. [209] introduced SCKansformer, a hybrid architecture tailored for fine-grained classification of bone marrow cells in microscopic images. Central to the model is a Kansformer encoder, which replaces the traditional MLP layers in Transformer blocks with KAN components to improve nonlinear feature extraction and model interpretability. This design is complemented by a spatial-channel convolutional encoder and a global-local attention encoder, together enhancing both redundancy reduction and context sensitivity in microimage analysis. The model was validated on a newly constructed Bone Marrow Cell Fine-Grained Classification Dataset, as well as publicly available benchmarks, demonstrating consistently superior accuracy and robustness across class-imbalanced and high-variability image sets. Ablation studies confirmed the critical contribution of the KAN-based module within the overall architecture, highlighting its value in handling morphological subtleties and long-tailed class distributions typical of clinical image data. These results together show the versatility of KANs in image classification tasks.
@@
In the domain of image segmentation, Li et al. [84] proposed U-KAN, a novel U-shaped architecture that integrates KANs into the encoder-decoder backbone of traditional convolutional segmentation models. Designed to address the challenges of low-contrast and high-variability tissue boundaries in medical imaging, U-KAN replaces MLP blocks with spline-parameterized KAN modules at the bottleneck and skip-connection layers. This hybrid structure enhances local adaptability and long-range dependency modeling without inflating parameter count. Applied to liver and tumor segmentation tasks from the LiTS dataset, U-KAN achieved state-of-the-art performance, surpassing Swin-UNet and TransUNet baselines in both Dice and HD95 metrics. As shown in Fig. 4(d), the model preserves sharper boundary details and finer anatomical structures, particularly in complex tumor regions. These results illustrate how KAN-enhanced modules can significantly improve segmentation quality in clinically relevant contexts, reinforcing their potential in data-driven biomedical image analysis.

### v2

@@ 4.1 Architecture and Training Strategies
To achieve this objective, the network parameters are iteratively updated using backpropagation wright2022deep , dampfhoffer2023backpropagation , ji2024comprehensive ; a procedure that computes the gradients of the loss function with respect to all trainable parameters. These gradients provide the necessary direction for optimization algorithms to adjust the network parameters toward minimizing the loss. Commonly used first-order optimization methods include stochastic gradient descent (SGD) newton2018stochastic , tian2023recent and the Adam optimizer kingma2014adam , both of which rely on gradient information to efficiently update the model during training. In contrast, second-order or quasi-Newton methods, such as the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm nawi2006improved and its limited-memory variant (L-BFGS) saputro2017limited , utilize curvature information to accelerate convergence. More recent advancements, such as the Self-Scaled Broyden (SSB) algorithms al2014broyden , kiyani2025optimizer , further enhance these quasi-Newton approaches by employing self-scaling strategies. These quasi-Newton methods iteratively construct approximations to the inverse Hessian matrix, allowing for more informed parameter updates by incorporating curvature information. This typically results in faster and more stable convergence compared to first-order methods, especially in high-dimensional or ill-conditioned optimization problems wright2006numerical , rafati2020quasi . To address the optimization challenges in data-driven KANs, it is worth considering hybrid strategies that combine the strengths of different optimizers. For instance, combining adaptive first-order methods like Adam kingma2014adam , which are effective for fast initial convergence, with second-order methods such as L-BFGS liu1989limited , saputro2017limited , which offer precise updates near local optima, may significantly improve training efficiency. This approach can be particularly beneficial for KANs, where the optimization involves a large number of parameters and the joint tuning of both the shape and location of spline-based transformations de1978practical , bohra2020learning , fakhoury2022exsplinet . In this context, Adam can help navigate the early stages of training with noisy gradients, while L-BFGS can refine the solution more effectively in the later stages.
@@
Recent advances in the design of data-driven KAN architectures have centered on structural innovations aimed at improving expressivity, interpretability, and efficiency. One direction involves hybrid basis function integration, as demonstrated in BSRBF-KAN, which combines B-splines with radial basis functions to enhance local approximation accuracy ta2024bsrbf . This hybrid formulation is implemented through a combined transformation layer applied to normalized inputs. The authors highlighted that, while both basis types contribute to performance, ablation studies suggest that B-splines play a more dominant role in preserving generalization. Additionally, Ta ta2024bsrbf provided insights into model stability and convergence behavior, indicating BSRBF-KAN’s potential as a flexible template for further exploration of basis function combinations in KANs. Other works, such as FC-KAN ta2024fc , propose function composition strategies by combining multiple elementary functions, such as wavelets, B-splines, and Gaussians, using element-wise operations or polynomial mappings to boost learning capacity. These methods show improved accuracy at the cost of increased training time, highlighting a trade-off between representational richness and computational load. Another important line of work focuses on initialization strategies for spline-based KANs. Rigas et al. rigas2025initialization systematically evaluated both theory-driven schemes, inspired by LeCun lecun2002efficient and Glorot glorot2010understanding initializations, and an empirical family of power-law strategies, showing that initialization plays a decisive role in convergence speed and predictive accuracy across regression and PDE benchmarks. In a different direction, Polar & Poluektov polar2021probabilistic proposed a probabilistic extension of the KAN by combining it with the divisive data re-sorting (DDR) method. While conventional regression models focus on expected values or confidence intervals, their method enables KANs to represent full probability distributions of outputs under aleatoric uncertainty, efficiently capturing input-dependent and multimodal behaviors with relatively low computational cost. Moreover, Mostajeran & Faroughi mostajeran2024epi proposed physics-guided parallel and serial KAN architectures with Chebyshev basis functions to evaluate their effectiveness in accurately modeling complex datasets that represent elasto-plastic material behavior. These structural adaptations reflect the growing interest in tailoring KAN frameworks to specific task demands, offering pathways toward more flexible and application-aware neural architectures.
@@
Alongside structural enrichment, another key research direction involves simplifying data-driven KAN architectures to reduce computational overhead without compromising performance. The HKAN model dudek2025hkan introduces a novel hierarchical design that replaces iterative backpropagation with a non-iterative, layer-wise least squares fitting procedure, eliminating gradient computation. This architecture adopts a stacking mechanism in which each layer incrementally refines the previous output, enhancing model stability and interpretability while remaining computationally efficient. Meanwhile, the PRKAN architecture ta2025prkan targets parameter reduction by combining convolutional layers, pooling, attention mechanisms, and dimension summation strategies to drastically lower the number of learnable parameters. Despite reduced complexity, PRKAN maintains competitive performance in image classification, leveraging normalization and hybrid basis functions like GRBFs. In parallel, the Multifidelity KAN (MFKAN) framework howard2025multifidelity provides a method for combining low- and high-fidelity datasets, thereby reducing the need for costly high-fidelity data. By using both linear and nonlinear relationships between datasets, MFKANs become more reliable when data is limited. This reflects a shift toward lightweight, scalable KAN architectures, balancing structural simplicity with functional robustness. Warin warin2024p1 also introduced the P1-KAN architecture, which employs compact piecewise-linear basis functions with bounded support. This formulation simplifies the representation while maintaining high approximation accuracy, particularly for irregular or non-smooth functions in high-dimensional spaces.
@@ 4.2 Applications
The adoption of data-driven KANs has been on the rise for a diverse array of tasks, including forecasting, regression, classification, and segmentation applications in both structured tabular domains and high-dimensional vision tasks. These models leverage the flexible function approximation capabilities to learn complex mappings directly from high-dimensional data. A summary of recent application-oriented advances in data-driven KANs is provided in Table 2.
@@
EPi-cKAN Developed physics-guided serial and parallel architectures to model nonlinear elasto-plastic behavior of granular materials; Used Chebyshev polynomials as basis functions; Applied to predict stress–strain responses. mostajeran2024epi
@@
Time-Series KAN Developed a KAN framework for time series forecasting; Used B-spline basis functions (degree 3 and grid size 5); Applied to forecast real-world satellite traffic using normalized hourly data from the 5G-STARDUST project. vaca2024kolmogorov
@@
Geotechnical KAN Developed a KAN model to predict settlement and residual strength after liquefaction; Used B-spline basis functions; Applied to field and lab data from case studies, including risk assessment in Patna and railway embankment analysis. karakacs2024novel
@@
COEFF-KANs Developed a two-stage modeling strategy to predict Coulombic Efficiency (CE) of lithium-metal battery electrolytes; Applied to a real-world Li/Cu half-cell dataset to predict log-transformed CE values. li2024coeff
@@
SCKansformer Developed a fine-grained classification model to improve diagnostic accuracy in hematological malignancies by addressing limitations of existing microimage analysis methods; Applied to over 10,000 cell images across three datasets (BMCD-FGCD, PBC, ALL-IDB). chen2024sckansformer
@@
U-KAN Developed a KAN-enhanced convolutional architecture to improve nonlinear representation and interpretability in image segmentation and generative tasks; Used tokenized KAN layers; Applied to diverse imaging datasets (BUSI, GlaS, and others). li2025u
@@
KACQ-DCNN Developed a classical–quantum hybrid neural architecture; Used B-spline basis functions within a dual-channel design to model nonlinear activation patterns and capture complex feature relationships; Applied to structured data for binary classification. jahin2024kacq
KAN-IDIR Developed an implicit neural representation framework for deformable image registration with KANs; Utilized adapted Chebyshev KANs ; Applied to three medical imaging benchmarks: DIR-Lab lung CT, OASIS-1 brain MRI, ACDC cardiac MRI. drozdov2025learning
@@
In the context of regression tasks, Vaca-Rubio et al. vaca2024kolmogorov proposed a data-driven KAN architecture as a parameter-efficient and interpretable alternative to traditional MLPs for time series forecasting, with a specific focus on satellite traffic prediction. Inspired by KART, their model replaces conventional weight-activation pairs with spline-parametrized univariate transformations, enabling greater modeling flexibility. As shown in Fig. 4(a), the KAN model closely follows the true traffic trajectory, particularly during rapid fluctuations. This demonstrates KAN’s ability to capture high-frequency temporal features and long-range dependencies. Further analysis demonstrates the impact of KAN-specific hyperparameters, such as node count and grid size, in balancing model accuracy and computational cost. These findings establish KANs as a promising, resource-efficient approach for complex time series forecasting tasks. Beyond temporal forecasting, KANs have also been successfully applied to structured regression problems. For example, Karakaş karakacs2024novel employed a data-driven KAN model to predict liquefaction-induced settlements using a dataset derived from the Pohang earthquake region in South Korea, previously compiled by Park et al. park2020simple . The input features included depth, unit weight, cyclic stress ratio, and corrected SPT blow count, with settlement magnitude as the target variable. The study demonstrated that KANs effectively capture the nonlinear dependencies between geotechnical parameters and settlement outcomes, achieving high predictive accuracy across standard evaluation metrics such as $R^{2}$, Mean Absolute Error (MAE), and Root Mean Squared Error (RMSE). Notably, the model identified cyclic stress ratio and SPT blow count as the most influential predictors, aligning well with domain-specific understanding. These results emphasize the potential of KANs as reliable and interpretable tools for modeling complex phenomena in data-limited environments. Further extending the applicability of KANs to granular material modeling, Mostajeran & Faroughi mostajeran2024epi developed Chebyshev-based KAN variants to predict nonlinear elasto-plastic behavior under complex loading. Using parallel and serial KAN architectures, the models achieved accurate predictions of plastic shear strain $\gamma^{p}$ and mean effective stress $p$, even under unseen undrained loading paths such as $\xi=-4$, as shown in Fig. 4(b). These results demonstrate the effectiveness of KANs in capturing stress–strain responses in elasto-plastic materials with compact and interpretable architectures. Further broadening the landscape of KAN applications, Li et al. li2024coeff introduced COEFF-KAN, a two-stage framework leveraging Kolmogorov–Arnold Networks to predict Coulombic Efficiency (CE) in lithium battery electrolytes. In this structured regression task, the model inputs consisted of SMILES-based molecular embeddings weighted by their molar ratios, derived from a pre-trained MoLFormer transformer. These formulation-level descriptors were then passed to a downstream KAN model, enabling the prediction of logarithmic CE (LCE) with state-of-the-art accuracy. As depicted in Fig. 4(c), the model exhibits robust generalization, with parity plots (top) demonstrating tight alignment between predicted and actual LCE values across held-out samples, and ablation studies (bottom) confirming KAN’s superior performance. The study underscores KAN’s strengths in capturing complex, high-dimensional structure–property relationships from limited chemical data, while preserving interpretability through its modular, spline-based transformations.
@@
Figure 4: (a) KAN with learnable B-spline activations for time series forecasting. The forecast is based on real GEO satellite traffic data (hourly resolution, 168h context / 24h prediction) provided by the 5G-STARDUST project, highlighting KAN’s applicability to AI-driven satellite resource management. This panel is adopted from vaca2024kolmogorov . (b) Prediction of plastic shear strain $\gamma^{p}$ and mean effective stress $p$ using parallel and serial KAN architectures under an undrained loading path with $\xi=-4$, $p_{\text{in}}=375$ kPa, and $e_{\text{in}}=0.64$. The results are reproduced based on the method presented in mostajeran2024epi , along with the ground truth obtained from numerical integration. (c) Application of COEFF-KAN to electrolyte property prediction. Top: Parity plot comparing predicted and ground-truth values of logarithmic Coulombic Efficiency (LCE) across test samples, illustrating strong agreement and model generalization. Bottom: RMSE performance of KAN versus MLP across varying network depths and widths. This panel is reproduced from li2024coeff . (d) Channel-wise activation maps for explainability evaluation in tumor segmentation. Integrating KAN layers enhances spatial alignment between model attention and ground truth, improving plausibility in localization tasks. This panel is adopted from li2025u .
@@
In the context of classification tasks, Ta ta2024bsrbf introduced BSRBF-KAN to evaluate two standard classification benchmarks, MNIST and Fashion-MNIST, using a consistent network structure across all experiments. BSRBF-KAN achieved stable and competitive performance, reporting average accuracies of 97.55% on MNIST and 89.33% on Fashion-MNIST, while also exhibiting faster convergence and smoother loss decay across training epochs. This study revealed the importance of architectural components such as layer normalization and base output terms, while highlighting that B-spline components contributed more significantly to performance than RBFs. These results emphasize the potential of hybrid basis-function strategies within the KAN framework for enhancing expressivity and robustness in classification tasks. In the same line of KAN-based classification work, Chen et al. chen2024sckansformer introduced SCKansformer, a hybrid architecture tailored for fine-grained classification of bone marrow cells in microscopic images. Central to the model is a Kansformer encoder, which replaces the traditional MLP layers in Transformer blocks with KAN components to improve nonlinear feature extraction and model interpretability. This design is complemented by a spatial-channel convolutional encoder and a global-local attention encoder, together enhancing both redundancy reduction and context sensitivity in microimage analysis. The model was validated on a newly constructed Bone Marrow Cell Fine-Grained Classification Dataset, as well as publicly available benchmarks, demonstrating consistently superior accuracy and robustness across class-imbalanced and high-variability image sets. Ablation studies confirmed the critical contribution of the KAN-based module within the overall architecture, highlighting its value in handling morphological subtleties and long-tailed class distributions typical of clinical image data. These results together show the versatility of KANs in image classification tasks.
@@
In the domain of image segmentation, Li et al. li2025u proposed U-KAN, a novel U-shaped architecture that integrates KANs into the encoder-decoder backbone of traditional convolutional segmentation models. Designed to address the challenges of low-contrast and high-variability tissue boundaries in medical imaging, U-KAN replaces MLP blocks with spline-parameterized KAN modules at the bottleneck and skip-connection layers. This hybrid structure enhances local adaptability and long-range dependency modeling without inflating parameter count. Applied to liver and tumor segmentation tasks from the LiTS dataset, U-KAN achieved state-of-the-art performance, surpassing Swin-UNet and TransUNet baselines in both Dice and HD95 metrics. As shown in Fig. 4(d), the model preserves sharper boundary details and finer anatomical structures, particularly in complex tumor regions. These results illustrate how KAN-enhanced modules can significantly improve segmentation quality in clinically relevant contexts, reinforcing their potential in data-driven biomedical image analysis.